OpenStudy (anonymous):

12x = 3 - 2y 2x + 4y = 6

OpenStudy (anonymous):

@hartnn?

hartnn (hartnn):

substitution again ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

first , as simplification., divide 2nd equation by 2.

OpenStudy (anonymous):

Wait i messed up. /.\ the first equation is x = 3 - 2y.

hartnn (hartnn):

then, you can directly substitute that x in 2nd equation...

OpenStudy (anonymous):

2( 3-2y) + 4y = 6 ?

hartnn (hartnn):

yes, go on, solve further...if you make an error, i'll stop you and tell u.

OpenStudy (anonymous):

I got 6 - 8y = 6, but then i added 6 and then it went to 8y = 12 .. :/

OpenStudy (anonymous):

Hello? @hartnn?

hartnn (hartnn):

you get 6-4y+4y=6 that is 6=6 which is an identity, hence there will be infinite solutions to this system of equation. sorry for late reply.

hartnn (hartnn):

sorry, you have to make a new question...

OpenStudy (anonymous):

But how do you get 6 = 6 ?

hartnn (hartnn):

\[2( 3-2y) + 4y = 6 \\ 6-4y+4y=6 \\ 6-0=6 \\ 6=6 \\identity\]

OpenStudy (anonymous):

Mkaii. :3 i need help on 1 more then im done.

hartnn (hartnn):

sure...